OpenStudy (anonymous):

Solutions for y=x^2 y=x

OpenStudy (hba):

Well lets see it like this, Subtract the two equations :)

OpenStudy (anonymous):

I prefer substitution, but both methods work

OpenStudy (hba):

Well you should go with substitution then @FlyinSolo_424

OpenStudy (anonymous):

y = x^2 y = x x = x^2 0 = x^2 - x 0 = x(x-1) x = 0,1

OpenStudy (hba):

Looks to me like your'e giving direct answer.

OpenStudy (hba):

Let her do the work. It is against the CoC.

OpenStudy (anonymous):

Elimination y = x^2 - (0)x y = (0)x^2 + x subtracting we have 0 = x^2 - x 0 = x(x-1) x=0,1

OpenStudy (anonymous):

@hba I am aware of the code of conduct

OpenStudy (hba):

@ChmE As a rule of thumb, we would prefer the askers to participate in the solution process instead of having the answerers doing all the work. thats not really our concern; we are aimed at being a study site, not a free answering service. We would rather promote learning the material as opposed to just handing over answers. This social contract is aimed at curbing people from using the site to cheat with ...

OpenStudy (hba):

@ChmE Well i saw what you did above -_-.

OpenStudy (callisto):

Taking a test?

OpenStudy (anonymous):

I wasnt saying that @hba ? I just need step by step because you guys are wanting me to do it on my own and I have no clue where to start nor finish. I didnt ask for the nswer.. i came on here for a step by step. i could go to yahoo or wolfram if I really wanted to..

OpenStudy (anonymous):

wanted a 'straight forward' anwser/

Parth (parthkohli):

Please, do not give answers to tests hereof.

OpenStudy (anonymous):

Steps to solving Systems of Equations: 1) Decide to use elimination or substitution (or matrices) 2) Elimination: If the equation has a common term this is normally faster. Common term meaning same coefficient already. You can also multiply an equation by a scalar to make two equations get a common term. a) once this term is eliminated by subtraction, you can solve for the remaining unknown variable through normal algebra steps 3) Substitution: This implies one equation is solved for an unknown variable and you substitute this into the same unknown in the other equation. This converts that equation into a single unknown equation and you can solve it for that unknown variable. This method is usually faster when an equation is already solved for one unknown. happy @hba ?!

OpenStudy (anonymous):

Assessment or whatever the heck you call it.. its 10 questions... look like test to me.

OpenStudy (hba):

@ChmE Well why would i be happy ? I just don't want to see you breaking the rules :) Btw,I am very thankful to you for understanding :)

OpenStudy (anonymous):

If you graph these.. are the vertex the solutions?

OpenStudy (hba):

@FlyinSolo_424 Lets use Elimination,It's easy :)

OpenStudy (hba):

@FlyinSolo_424 Just subract these two equations, y=x^2 y=x

OpenStudy (hba):

The question which comes in your mind is " Why should i subract the two equations ? " We are subtracting these two equations so we could eliminate 'y' and find 'x'.

OpenStudy (anonymous):

Would we just be left with x?

OpenStudy (hba):

Exactly,we would be left with the x variables. :)

OpenStudy (hba):

Did you subtract,what did you get ? @FlyinSolo_424

OpenStudy (anonymous):

x=x ?

OpenStudy (hba):

y=x^2 y=x - - ------- Now subtract it again,You did it wrong there.

OpenStudy (anonymous):

i mean y=x

OpenStudy (hba):

What is y-y = and x^2-x =

OpenStudy (hba):

@FlyinSolo_424 Well if you do not understand this,we can use another way ?